OpenStudy (anonymous):

Find all values of Theta in the interval 0 degrees < theta <360 degrees that satisfy the equation sin2Theta = sintheta

OpenStudy (anonymous):

\[\sin(2x)=2\sin(x)\cos(x) => 2\sin(x)\cos(x)=\sin(x) => \cos(x)=1/2\] This is true when x=60 degrees or 300 degrees